OpenStudy (anonymous):

The formula for degrees Celsius is: C = (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 4 degrees Celsius is.

OpenStudy (anonymous):

That is not the correct formula to convert Celsius to Fahrenheit. Is that really what your problem says??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Interesting... very interesting. I'm not really sure how you should answer a question that is badly flawed. This is sort of like saying, "the formula to convert yards Y to feet F is F = 4 * Y, solve for Y"

OpenStudy (anonymous):

Basically, you just have to get to the point where F = something. So, start by looking at the original equation: C = (5/9)*(F-32) Divide both sides by (5/9) Remember, when dividing fractions, switch the numerator and denominator, and then multiply by that fraction. C*(9/5) = F-32. Then add 32: C*(9/5) + 32 = F The 2nd part indicates that you plug a value for C to get the F value: [(6)*9/5] + 32 = F 10.8 + 32 = F 42.8 = F

OpenStudy (anonymous):

Objection: assumes formula not yet admitted into the discussion.

OpenStudy (anonymous):

Yes, you do have to get to F= something, but if you start with the wrong formula, you won't get the right answers. The real right answer is as @anas2000 says... first of all, realize the question is broken, look up the REAL formula, and then solve for F.